# NotaGen: Advancing Musicality in Symbolic Music Generation with Large Language Model Training Paradigms

### v3

@@ Abstract
We introduce NotaGen, a symbolic music generation model aiming to explore the potential of producing high-quality classical sheet music. Inspired by the success of Large Language Models (LLMs), NotaGen adopts pre-training, fine-tuning, and reinforcement learning paradigms (henceforth referred to as the LLM training paradigms). It is pre-trained on 1.6M pieces of music, and then fine-tuned on approximately 9K high-quality classical compositions conditioned on “period-composer-instrumentation” prompts. For reinforcement learning, we propose the CLaMP-DPO method, which further enhances generation quality and controllability without requiring human annotations or predefined rewards. Our experiments demonstrate the efficacy of CLaMP-DPO in symbolic music generation models with different architectures and encoding schemes. Furthermore, subjective A/B tests show that NotaGen outperforms baseline models against human compositions, greatly advancing musical aesthetics in symbolic music generation.
@@ 1 Introduction
The pursuit of musicality is a core objective in music generation research, as it fundamentally shapes how we perceive and experience musical compositions. Symbolic music abstracts music into discrete symbols such as notes and beats, with MIDI and sheet music being the two predominant formats. These formats can efficiently model melody, harmony, instrumentation, etc., all of which are crucial for musicality.
@@
In this work, we introduce NotaGen (Musical Notation Generation), a symbolic music generation model focused on classical sheet music. Compared to MIDI generation, not only does sheet music generation aim to produce artistically refined music, but it also emphasizes proper voice arrangement and notation to create well-formatted sheets for performance and analysis. Furthermore, the challenge of sheet music generation is exacerbated by the diverse instrumentation and rich musicality inherent in classical music. The success of LLMs has motivated us to apply the training paradigms to sheet music generation. NotaGen is pre-trained on a corpus of over 1.6M sheets, and fine-tuned on a collection of approximately 9K high-quality classical pieces from 152 composers with “period-composer-instrumentation” (e.g.“Baroque-Bach, Johann Sebastian-Keyboard”) prompts guiding conditional generation. During reinforcement learning, we introduce the CLaMP-DPO method to further optimize NotaGen’s musicality and controllability using the Direct Preference Optimization (DPO) Rafailov et al. (2024) algorithm. In this approach, CLaMP 2 Wu et al. (2024b), a multimodal symbolic music information retrieval model, assigns generated samples as “chosen” or “rejected” based on references from the fine-tuning dataset. Our contributions are two-fold:
@@ 3.3.1 Pre-training
We cleaned the data rigorously to ensure the quality: pieces with fewer than 8 bars were excluded, and deduplication was performed based on MD5 checksums and edit distances between sheet texts. We also preprocessed the text annotations, retaining music-related content such as tempo and expression hints, while removing irrelevant content like lyrics and background information. This resulted in a final pre-training dataset of 1.6M music sheets.
@@ Appendix C Prompt Set in Experiments
Continued on next page |
@@
Continued on next column |

### v4

@@ Abstract
We introduce NotaGen, a symbolic music generation model aiming to explore the potential of producing high-quality classical sheet music. Inspired by the success of Large Language Models (LLMs), NotaGen adopts pre-training, fine-tuning, and reinforcement learning paradigms (henceforth referred to as the LLM training paradigms). It is pre-trained on 1.6M pieces of music in ABC notation, and then fine-tuned on approximately 9K high-quality classical compositions conditioned on “period-composer-instrumentation” prompts. For reinforcement learning, we propose the CLaMP-DPO method, which further enhances generation quality and controllability without requiring human annotations or predefined rewards. Our experiments demonstrate the efficacy of CLaMP-DPO in symbolic music generation models with different architectures and encoding schemes. Furthermore, subjective A/B tests show that NotaGen outperforms baseline models against human compositions, greatly advancing musical aesthetics in symbolic music generation.
@@ 1 Introduction
The pursuit of musicality is a core objective in music generation research, as it fundamentally shapes how we perceive and experience musical compositions. Symbolic music abstracts music into discrete symbols such as notes and beats, with performance signals (i.e., MIDI) and sheet music (e.g., ABC notation, MusicXML) being the two predominant modalities. Both of them can efficiently model melody, harmony, instrumentation, etc., all of which are crucial for musicality.
@@
In this work, we introduce NotaGen (Musical Notation Generation), a symbolic music generation model focused on classical sheet music. Compared to MIDI generation, sheet music generation not only aims to produce artistically refined music, but also emphasizes proper voice arrangement and notation to create well-formatted sheets for performance and analysis. Furthermore, the challenge of sheet music generation is exacerbated by the diverse instrumentation and rich musicality inherent in classical music. The success of LLMs has motivated us to apply the training paradigms to sheet music generation. NotaGen is pre-trained on a corpus of over 1.6M sheets in ABC notation, and fine-tuned on a collection of approximately 9K high-quality classical pieces from 152 composers with “period-composer-instrumentation” (e.g.“Baroque-Bach, Johann Sebastian-Keyboard”) prompts guiding conditional generation. During reinforcement learning, we introduce the CLaMP-DPO method to further optimize NotaGen’s musicality and controllability using the Direct Preference Optimization (DPO) Rafailov et al. (2024) algorithm. In this approach, CLaMP 2 Wu et al. (2024b), a multimodal symbolic music information retrieval model, assigns generated samples as “chosen” or “rejected” based on references from the fine-tuning dataset. Our contributions are two-fold:
@@ 3.3.1 Pre-training
The pre-training stage utilized a carefully curated internal-use dataset comprising 1.6M ABC notation sheets. We also preprocessed the text annotations, retaining music-related content such as tempo and expression hints, while removing irrelevant content like lyrics and background information.
@@ Appendix C Prompt Set in Experiments
| Continued on next page | ||
@@
| Continued on next column | ||